Please solve the equation by factoring:
3x^2-24x+48=0

Answer by Nate(3500)

3x^2-24x+48=0 original
x^2-8x+16=0 divide by 3
(x-4)(x-4)=0 the sum of -4 and -4 to get -8; the product of -4 and -4 to get 16
x-4=0; x-4=0
x=4; x=4
the answer is 4